Using categories of price electricity, how would you advise a companies Board Director or employers on the effects of price changes on quantity demanded and total revenue of product.

( Hint: P*Q=TR. Where P=Price, Q=Quentity, TR=Total Revenue )
Answer #1

Ans) Price elasticity of demand is the responsiveness of quantity demanded to change in price.

If even the small price change brings a significant change in quantity demanded, good is said to be elastic. In this case, increasing price is not recommended as people can easily alter their demand and it will reduce the revenue.

If even the significant change in price does not change the quantity demanded much, then, good is said to be inelastic. In this case increasing price can bring more revenue as people cannot alter their demand easily.

Changing price in case of elastic good÷

# Increasing price from $10 to $12, changed quantity demanded from 30 to 15 units

Revenue when price was $10 = 10×30 = $300

Revenue when price is $12= 12×15 = $180

Changing price in case of inelastic good÷

# Increasing price from $10 to $15, changed quantity demanded from 30 to 25

Revenue when price was $10 = 10×30 = $300

Revenue when price is $15 = 15×25 = $375